# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v2

@@ Abstract
Performance prediction is a method to estimate the performance of Language Models (LMs) on various Natural Language Processing (NLP) tasks, mitigating computational costs associated with model capacity and data for fine-tuning. Our paper introduces ProxyLM, a scalable framework for predicting LM performance using proxy models in multilingual tasks. These proxy models act as surrogates, approximating the performance of the LM of interest. By leveraging proxy models, ProxyLM significantly reduces computational overhead on task evaluations, achieving up to a 37.08$\times$ speedup compared to traditional methods, even with our smallest proxy models. Additionally, our methodology showcases adaptability to previously unseen languages in pre-trained LMs, outperforming the state-of-the-art performance by 1.89$\times$ as measured by root-mean-square error (RMSE). This framework streamlines model selection, enabling efficient deployment and iterative LM enhancements without extensive computational resources.
@@ 1 Introduction
Language Models (LMs) have become increasingly valuable for assessing Natural Language Processing (NLP) tasks Raffel et al. (2020); Brown et al. (2020); Costa-jussà et al. (2022b); Touvron et al. (2023a, b); Le Scao et al. (2023). However, fine-tuning and evaluating these models are resource-intensive processes in terms of both computation and time. These costs escalate with model size, especially when experimenting across multiple datasets. As highlighted in Kaplan et al. (2020), there is a scaling law that applies to both model and dataset sizes, and computational demands, indicating that larger models and broader datasets require increased computational resources. Modeling low-resource languages (LRLs) in multilingual contexts presents a range of challenges. One significant challenge is the limited data availability, which hampers effective fine-tuning processes Gu et al. (2018); Adilazuarda et al. (2024), making model adaptation through fine-tuning a challenging task Zoph et al. (2016); Liu et al. (2021). Another critical issue is the lack of pre-training data for numerous regional languages, such as Southeast Asian languages Winata et al. (2022, 2024); Yong et al. (2024), with many languages being omitted during the pre-training phase of multilingual LMs.
@@
Given the limited academic computational resources for LM fine-tuning and inadequate datasets for LRLs, performance prediction is an efficient method to alleviate the dependency on extensive resources by leveraging past performance records on an NLP task. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008a); Srinivasan et al. (2021); Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022a) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART.
In this paper, we propose ProxyLM,  [^1] We release our code at https://github.com/davidanugraha/proxylm a framework to predict LMs performance by utilizing proxy models on LRLs. Proxy models are defined as substitute models, wherein the performance of these substitute models are used to estimate the performance of another LM. This other model can be significantly larger than our proxy models. For optimizing the prediction, we utilize much smaller LMs as proxy models and off-the-shelf models without further tuning. This approach is very scalable to multiple proxy models and task-agnostic to any modalities, thus it can be applied to any downstream tasks. This study focuses on machine translation tasks and our approach outperforms the existing work from Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), which opens a new avenue to employ LMs for model performance prediction. Therefore the contribution of our paper can be summarized in three-fold:
@@
    We introduce ProxyLM, an efficient and scalable framework designed to predict the performance of LMs. This framework significantly reduces the computational costs associated with fine-tuning and inference during model selection.
@@
    We demonstrate the effectiveness and robustness of ProxyLM across 21 dataset sources and 50 languages on two estimated LM architectures. Our framework substantially outperforms all existing baselines in English-centric, many-to-many languages, and cross-dataset settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs, surpassing the state-of-the-art performance measured with root-mean-square error (RMSE) by 1.89$\times$.
@@
    We also provide a time analysis comparing the fine-tuning duration of proxy models to direct model fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a 37.08$\times$ speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
@@ 2.1 ProxyLM
We can approximate $f$ by transforming Equation 1 into a regression task with a regressor function $g$, which will be trained on past performance records. The regressor takes dataset features $\Phi(\mathcal{D},\mathcal{D}^{\prime})$ to identify the characteristics of the training and test datasets, as well as the distribution shift between them. It also takes language features $\Psi(\mathcal{L}_{s},\mathcal{L}_{t})$ to measure the dissimilarities between the source and target languages. This can be formulated as follows:
@@
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model $\mathcal{M}$. Formally, let $\mathcal{M}_{p}=[\mathcal{M}_{p}^{1},\dots,\mathcal{M}_{p}^{N}]$ be a set of $N$ proxy models. To integrate the information from these proxy models, we modify Equation 2 as follows:
@@ 2.2 ProxyLM Features
Language Features
We use URIEL Typological Database Littell et al. (2017) similar to Xia et al. (2020a) including geographic, genetic, inventory, syntactic, phonological, and featural distance. The language features are useful to provide a language-specific representation to the regressor.
Dataset Features
We extract 6 features from the dataset, including train size, vocab size, average sentence length, word overlap, Type-Token Ratio (TTR), and TTR distance from $\mathcal{D}$ and $\mathcal{D^{\prime}}$ based on  Xia et al. (2020a). We will refer to these features and language features combined as NLPerf features. Furthermore, we incorporate the distribution shift information between the training and test datasets using Jensen-Shannon Divergence (JSD) as described by  Khiu et al. (2024). In addition, we include term frequency-inverse document frequency (TF-IDF) and sentence similarity with Sentence-BERT  Reimers and Gurevych (2019).
Proxy Models Features
@@ 3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records are then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in A.3.
@@ 3.1 Datasets
We evaluate our approach using two types of datasets: English-centric and Many-to-Many Languages. The English-centric dataset involves English serving as either the source or target language. Our English-centric dataset comes from the MT560 Gowda et al. (2021) dataset, where we curated 19 datasets and selected 44 languages out of 500 for evaluation. Furthermore, we use the FLoRes-200 dataset Costa-jussà et al. (2022b) for additional validation and test sets. In contrast, the Many-to-Many Languages dataset allows any language to act as the source or target. We use the NusaTranslation dataset  Cahyawijaya et al. (2023) as our Many-to-Many Languages dataset, which comprises parallel texts in 12 Indonesian regional languages. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. Both datasets encompass 50 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in Section A.1 in the Appendix.
3.2 Models
Estimated LMs
@@ 3.2 Models
We employ two estimated LMs: M2M100 1.2B Fan et al. (2021) and NLLB 1.3B Costa-jussà et al. (2022b). Each estimated model is fine-tuned using a standard next token prediction objective on the training set.
Proxy Models
@@
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB  Costa-jussà et al. (2022b). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. Model details are provided in Section  A.2 in the Appendix. The evaluation is done using SentencePiece BLEU (spBLEU) Goyal et al. (2022), as it has been demonstrated to be a fair metric in multilingual settings, particularly in low-resource settings. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs.
Regressors
@@
We utilize XGBoost Chen and Guestrin (2016), LGBM Ye et al. (2021a), Poly2 Khiu et al. (2024), and Poly3 Khiu et al. (2024) as our regressors. In most of our experiments, we apply XGBoost as our default regressor because we find it to be the best-performing model, while the other regressors serve as baselines. Specifically for the Many-to-Many Languages setting, Matrix Factorization with context features (MF) is used as an additional baseline Schram et al. (2023). We do not apply MF to our English-centric setting because MF requires the performance records to be structured in two dimensions—one for the source language and one for the target language. In the English-centric setting, this would result in a sparse matrix with only one fully populated row or column, corresponding to English, making MF impractical for this setup.
Models Random LOLO Unseen
@@
NLPerf Xia et al. (2020a) with different regressors
XGBoost 8.09 $\pm$ 0.44 8.40 $\pm$ 0.41 8.81 $\pm$ 0.39 12.35 $\pm$ 0.52 8.22 $\pm$ 0.46 9.17
Poly2 Khiu et al. (2024) 10.00 $\pm$ 0.46 12.30 $\pm$ 0.48 10.54 $\pm$ 0.00 13.28 $\pm$ 0.00 9.91 $\pm$ 0.00 11.21
Poly3 Khiu et al. (2024) 9.93 $\pm$ 0.52 12.33 $\pm$ 0.46 10.66 $\pm$ 0.00 14.06 $\pm$ 0.00 10.57 $\pm$ 0.00 11.51
LGBM Ye et al. (2021a) 8.18 $\pm$ 0.34 8.59 $\pm$ 0.33 9.00 $\pm$ 0.49 12.41 $\pm$ 0.53 8.59 $\pm$ 0.69 9.35
@@
Transformer 5.29 $\pm$ 0.38 7.65 $\pm$ 0.47 6.25 $\pm$ 0.29 11.33 $\pm$ 0.50 6.54 $\pm$ 0.30 7.41
SMaLL-100 4.26 $\pm$ 0.35 7.02 $\pm$ 0.44 4.33 $\pm$ 0.24 9.70 $\pm$ 0.37 4.54 $\pm$ 0.41 5.97
SMaLL-100 (No FT) 5.93 $\pm$ 0.23 6.73 $\pm$ 0.49 6.33 $\pm$ 0.28 10.66 $\pm$ 0.44 5.77 $\pm$ 6.06 7.08
Estimated Model (No FT) 5.76 $\pm$ 0.19 4.29 $\pm$ 0.31 6.23 $\pm$ 0.34 4.59 $\pm$ 0.22 5.67 $\pm$ 0.41 5.31
Ensemble 3.79 $\pm$ 0.30 3.79 $\pm$ 0.25 3.90 $\pm$ 0.22 4.14 $\pm$ 0.24 4.36 $\pm$ 0.22 4.00
Table 1: English-centric test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result can be seen in Section B in the Appendix.
Models M2M100 NLLB
LRL MRL LRL MRL
Transformer 6.14 6.67 11.60 10.26
SMaLL-100 4.28 4.55 10.18 7.83
SMaLL-100 (No FT) 6.17 6.96 7.90 11.38
Estimated Model (No FT) 6.15 6.57 4.66 4.32
Ensemble 3.93 3.80 4.10 4.15
Table 2: English-centric LOLO test results in average RMSE by language vitality (lower is better). “No FT" denotes “no fine-tuning". “LRL" and “MRL" denote low-resource languages and medium-resource languages, based on Joshi et al. (2021). The full mapping is in Table 8 & 9 in Section A.1 in the Appendix.
Models Random LOLO
[…]
NLPerf Xia et al. (2020a) with different regressors
XGBoost 2.45 $\pm$ 0.30 1.21 $\pm$ 0.29 7.83 $\pm$ 0.23 8.28 $\pm$ 0.31 4.94
LGBM  Ye et al. (2021a) 2.66 $\pm$ 0.22 1.76 $\pm$ 0.29 7.91 $\pm$ 0.01 8.06 $\pm$ 0.00 5.10
ProxyLM (Ours) with different proxy models
Transformer 2.56 $\pm$ 0.43 1.70 $\pm$ 0.20 5.65 $\pm$ 0.23 6.24 $\pm$ 0.34 4.04
SMaLL-100 2.56 $\pm$ 0.33 1.65 $\pm$ 0.44 4.85 $\pm$ 0.36 5.14 $\pm$ 0.46 3.55
Estimated Model (No FT) 2.38 $\pm$ 0.36 1.27 $\pm$ 0.30 5.10 $\pm$ 0.28 5.50 $\pm$ 0.26 3.56
Ensemble 2.41 $\pm$ 0.28 1.56 $\pm$ 0.35 3.73 $\pm$ 0.23 3.79 $\pm$ 0.19 2.90
Table 3: Many-to-Many Languages test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". We do not cover the Unseen setting as most of the languages are not covered by both M2M100 and NLLB in the Many-to-Many Languages dataset. Avg represents the average of the results across the row. The reported results are experiments using XGBoost regressor. Ensemble denotes combining all four proxy models.

### v3

@@ Abstract
Performance prediction is a method to estimate the performance of Language Models (LMs) on various Natural Language Processing (NLP) tasks, mitigating computational costs associated with model capacity and data for fine-tuning. Our paper presents ProxyLM, a scalable task- and language-agnostic framework designed to predict the performance of LMs using proxy models. These proxy models act as surrogates, approximating the performance of the LM of interest. By leveraging these proxy models, ProxyLM significantly reduces computational overhead in task evaluations, achieving up to a 37.08$\times$ speedup over traditional methods, even with our smallest proxy models. Our results across multiple multilingual NLP tasks and various robustness tests demonstrate that ProxyLM not only adapts well to previously unseen languages in pre-trained LMs, but also generalizes effectively across different datasets, outperforming the state-of-the-art by at least 1.78$\times$ in terms of root-mean-square error (RMSE).
ProxyLM: Predicting Language Model Performance on Multilingual
Tasks via Proxy Models
David Anugraha1, Genta Indra Winata2thanks: The work was conducted outside Capital One., Chenyue Li1, Patrick Amadeus Irawan3, En-Shiun Annie Lee1,4 1University of Toronto  2Capital One 3Institut Teknologi Bandung  4Ontario Tech University anugraha@cs.toronto.edu, genta.winata@capitalone.com, {jimschenchen,patrickamadeusirawan}@gmail.com, annie.lee@ontariotechu.ca
@@ 1 Introduction
Language Models (LMs) have become increasingly valuable for assessing Natural Language Processing (NLP) tasks Raffel et al. (2020); Brown et al. (2020); Costa-jussà et al. (2022); Touvron et al. (2023a, b); Workshop et al. (2022). However, fine-tuning and evaluating these models are resource-intensive processes in terms of both computation and time. These costs escalate with model size, especially when experimenting across multiple datasets. As highlighted in Kaplan et al. (2020), there is a scaling law that applies to both model and dataset sizes, and computational demands, indicating that larger models and broader datasets require increased computational resources. Modeling low-resource languages (LRLs) in multilingual contexts presents a range of challenges. One significant challenge is the limited data availability, which hampers effective fine-tuning processes Gu et al. (2018); Adilazuarda et al. (2024), making model adaptation through fine-tuning a challenging task Zoph et al. (2016); Liu et al. (2021). Another critical issue is the lack of pre-training data for numerous regional languages, such as Southeast Asian languages Winata et al. (2022, 2024b); Yong et al. (2024), with many languages being omitted during the pre-training phase of multilingual LMs.
@@
Given the limited academic computational resources for LM fine-tuning and inadequate datasets for LRLs, performance prediction is an efficient method to alleviate the dependency on extensive resources by leveraging past performance records on an NLP task. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008); Srinivasan et al. (2021); Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART and other LMs Zhu et al. (2023).
In this paper, we propose ProxyLM, [^1] We release our code at https://github.com/davidanugraha/proxylm. a scalable task- and language- agnostic framework to predict LM performance by utilizing proxy models. Proxy models are defined as substitute models, wherein the performance of these substitute models are used to estimate the performance of another LM. This other model can be significantly larger than our proxy models. For optimizing the prediction, we utilize much smaller LMs as proxy models and off-the-shelf models without further tuning. This approach is very scalable to multiple proxy models and task-agnostic to any modalities, thus it can be applied to any downstream tasks. This study focuses on three multilingual tasks, covering machine translation (MT), intent classification, and slot filling. Our approach outperforms the existing work from Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024), which opens a new avenue to employ LMs for model performance prediction. Therefore the contribution of our paper can be summarized in three-fold:
@@
    We introduce ProxyLM, an efficient and scalable task- and language-agnostic framework designed to predict the performance of LMs. This framework significantly reduces the computational costs associated with fine-tuning and inference during model selection.
@@
    We demonstrate the effectiveness and robustness of ProxyLM across 34 dataset sources and 56 languages for MT, and 51 languages for intent classification and slot filling, each on two estimated LMs. Our framework substantially outperforms all existing baselines in all NLP tasks, datasets, and settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs and across different datasets, surpassing the state-of-the-art performance measured with root-mean-square error (RMSE) by at least 1.78$\times$.
@@
    We also provide a time analysis comparing the deployment of proxy models for performance prediction to direct LM fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a 37.08$\times$ speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
@@ 2.1 ProxyLM
We can approximate $f$ by transforming Equation 1 into a regression task with a regressor function $g$, which will be trained on past performance records. Previous works  Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024) formulate regressor that takes dataset features $\Phi(\mathcal{D},\mathcal{D}^{\prime})$ to identify the characteristics of the training and test datasets, as well as the distribution shift between them. It also takes language features $\Psi(\mathcal{L}_{s},\mathcal{L}_{t})$ to measure the similarities between the source and target languages. This can be formulated as follows:
@@
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model $\mathcal{M}$, which addresses the gap in previous works for not accounting $\mathcal{M}$. Formally, let $\mathcal{M}_{p}=[\mathcal{M}_{p}^{1},\dots,\mathcal{M}_{p}^{N}]$ be a set of $N$ proxy models. To integrate the information from these proxy models, we modify Equation 2 as follows:
@@ 2.2 ProxyLM Features
Language Features.
We use URIEL Typological Database Littell et al. (2017) similar to Xia et al. (2020) including geographic, genetic, inventory, syntactic, phonological, and featural distance. The language features are useful to provide a language-specific representation to the regressor.
Dataset Features.
We extract six features from the dataset, including train size, vocab size, average sentence length, word overlap, Type-Token Ratio (TTR), and TTR distance from $\mathcal{D}$ and $\mathcal{D^{\prime}}$ based on  Xia et al. (2020). We will refer to these features and language features combined as NLPerf features. Furthermore, we incorporate the distribution shift information between the training and test datasets using Jensen-Shannon Divergence (JSD) as described by Khiu et al. (2024). In addition, we include the cosine similarity of term frequency-inverse document frequency (TF-IDF) representations and sentence embeddings using Sentence-BERT Reimers and Gurevych (2019). Details on how these features are computed can be found in Appendix Section A.
Proxy Models Features.
@@ 3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records are then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in B.3.
@@ 3.1 Datasets
We use two types of datasets: English-centric and Many-to-Many Languages. The English-centric dataset involves English serving as either the source or target language. Our English-centric dataset comes from the MT560 Gowda et al. (2021) dataset, where we curate 32 datasets and select 50 languages out of 500 for evaluation. Furthermore, we use the FLoRes-200 dataset Costa-jussà et al. (2022) for additional validation and test sets. These datasets consist of translations with varying quality across diverse domains, presenting significant challenges for a robust performance prediction. In contrast, the Many-to-Many Languages dataset allows any language to act as the source or target. We use the NusaTranslation dataset Cahyawijaya et al. (2023) as our Many-to-Many Languages dataset, which comprises parallel texts in 12 Indonesian regional languages. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. Both datasets encompass 56 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in the Appendix Section B.1.
Intent Classification and Slot Filling.
We use MASSIVE FitzGerald et al. (2022) as our dataset encompassing 51 languages. We fine-tune each LM on all languages and evaluate it on all languages. Detailed language insights are available in the Appendix Section B.1.
3.2 Estimated LMs
Machine Translation.
@@ 3.2 Estimated LMs
We employ two estimated LMs: M2M100 1.2B Fan et al. (2021) and NLLB 1.3B Costa-jussà et al. (2022). Each estimated model is fine-tuned using a standard next-token prediction objective on the training set.
Intent Classification and Slot Filling.
We employ two decoder-only estimated LMs: LLaMA-3 Instruct (8B)  Dubey et al. (2024) and Aya-23 (8B) Aryabumi et al. (2024). Both models are fine-tuned using supervised fine-tuning (SFT) combined with Low-Rank Adaptation (LoRA) Hu et al. (2021) on the training set.
3.3 Proxy Models
Machine Translation.
@@ 3.3 Proxy Models
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB  Costa-jussà et al. (2022). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs. Model details are provided in the Appendix Section  B.2. The evaluation is primarily conducted using SentencePiece BLEU (spBLEU) Goyal et al. (2022), which has proven to be a reliable metric in multilingual and LRLs. For a more comprehensive assessment, we also use COMET-22 Rei et al. (2022), as it shows a high correlation with human judgments Freitag et al. (2023). However, COMET-22 is applied to only a subset of the English-centric dataset due to its limited language coverage, especially to LRLs.
Intent Classification and Slot Filling.
We utilize three decoder-only models: SmolLM (135M and 360M) Allal et al. (2024) and BLOOMZ (560M) Muennighoff et al. (2022). Model details are provided in the Appendix Section  B.2. The evaluation is done using accuracy for intent classification and micro-F1 for slot filling.
3.4 Regressor Models
@@
We utilize XGBoost Chen and Guestrin (2016), LGBM Ye et al. (2021), Poly2 Khiu et al. (2024), and Poly3 Khiu et al. (2024) as our regressors. In most of our experiments, we apply XGBoost as our default regressor because we find it to be the best-performing model based on the cross-validation during training, while the other regressors serve as baselines. Specifically for MT in the Many-to-Many Languages setting, Matrix Factorization with context features (MF) is used as an additional baseline Schram et al. (2023). We do not apply MF to our English-centric setting because MF requires the performance records to be structured in two dimensions—one for the source language and one for the target language. In the English-centric setting, this would result in a sparse matrix with only one fully populated row or column, corresponding to English, making MF impractical for this setup.
English-centric Many-to-Many
Models Random LOLO Avg. Random LOLO Avg.
@@
NLPerf Xia et al. (2020) with different regressors
XGBoost 7.69 7.73 9.20 12.92 9.39 2.45 1.11 7.83 8.28 4.94
Poly2 Khiu et al. (2024) 11.21 16.23 15.55 43.02 21.50 4.70 4.68 7.07 7.90 6.09
Poly3 Khiu et al. (2024) 11.00 15.64 62.29 236.29 81.31 4.60 4.64 7.26 8.01 6.13
LGBM Ye et al. (2021) 7.88 8.15 9.71 12.81 9.64 3.65 2.60 7.08 7.14 5.12
@@
Transformer 4.68 7.22 6.18 11.78 7.47 2.56 1.70 5.65 6.24 4.04
SMaLL-100 4.07 6.33 4.59 10.33 6.33 2.56 1.65 4.85 5.14 3.55
SMaLL-100 (No FT) 5.27 6.04 6.28 10.94 7.13 2.44 1.34 6.93 7.25 4.49
Estimated Model (No FT) 5.23 4.15 6.18 5.42 5.25 2.38 1.27 5.10 5.50 3.56
Ensemble 3.21 3.68 3.74 4.94 3.89 2.41 1.56 3.73 3.79 2.90
Table 1: MT test results on English-centric and Many-to-Many Languages datasets using spBLEU in average RMSE (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. Avg represents the average of the results across the row for each respective dataset. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.